If a typical tree absorbs 2000 kg of CO2 per day, then how much kg of CO2 will it absorb in one month (30 days) and in one year (365 days)....?


1
Expert's answer
2021-08-29T14:39:46-0400

For one day amount of CO2 absorbed = 2000kg

Let the number of days be ‘x’, therefore the amount of CO2 will be (2000x) kg

Within one month (30days), amount of CO2 absorbed = (2000×30) =60,000kg

Within one year (365days), amount of CO2 absorbed = (2000×365) kg=730,000kg